One number is 5 more than twice another number. The sum of the numbers is 35. Find the numbers.
Answer by bluemockingjay7(41)

1st number = x
2nd number = 2x + 5
sum = 35
x + 2x + 5 = 35
3x + 5 = 35
3x = 35 - 5
3x = 30
3x/3 = 30/3
x = 10
1st number = 10
2nd number = 2(10) + 5 = 20 + 5 = 25